Use the following table to answer the question below.

Expected ret. std. dev.
S&P500 13% 25%
ABC Fund 16% 27%
T-bill 4%
Borrow 6%

What is the highest fee that a client who is currently borrowing would be willing to pay to invest in your fund instead of S&P500?

Round your answer to 4 decimal places. For example, if your answer is 3.205%, then please write down 0.0321.
Answer #1

Market price of risk for S&P500 is ( 0.13 - .06) / 0.25 = 0.28

Market price of risk for ABC fund is (.16 - 0.06) / 0.27 = 0.37037

Highest fee that a client who is currently borrowing would be willing to pay to invest in ABC fund instead of S&P500 is 0.37037 - 0.28 = 0.0904

i.e. 9.04 %